(1 point) If C is the line segment from (7,4) to, (0,0), find the value of the line integral: Sc(8y2 i + 2x j). dr = 952/3
Answer #1

Į (8y? 13 + 2 m3).d? The line segment from (7,4) to (ao) is Y-Y, = 3 - 1 (0-1) =)y=4 = orang (x-1) 34-4= (-7) =) Y = 4 dy=4,